The equation y= -1777x + 27,153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is x=0 corresponds to 2000, x=3 corresponds to 2003, x=4 is 2004 and so on. Predict the number of gun deaths in the year 2004 and 2007 and predict in what year the number of gun deaths will be 11,160Predicted gun deaths in 2000 and 2004 would be?? And what year will there be 11160 gun deaths?
How do I figure this out???? I was doing so good now am stuck again.
Will someone please help with this word problem??

Answer by checkley77(12844)

y=-1,777x+27,153
for x=4
y=-1,777*4+27,153
y=-7,108+27,153
y=20,045 for the number in 2004.
for x=7
y=-1,777*7+27,153
y=-12,439+27,153
14,714 for the number in 2007.
11,160=-1,777x+27,153
1,777x=27,153-11,160
1,777x=15,993
x=15,993/1,777
x=9 or the year 2009 the projected deaths will be 11,160.